convert these repeating decimals .090909 .999 .909090
Answer by Alan3354(69443)

convert these repeating decimals
.090909
n = 0.090909...
100n = 9.090909...
- n = 0.090909...
99n = 9
n = 9/99
n = 1/11
-----------
.999
1n = 0.999...
10n = 9.999...
-1n = -.999
9n = 9
n = 1
------------
.909090
It's 10 times the 1st one,
= 10/11